# MROVSeg: Breaking the Resolution Curse of Vision-Language Models in Open-Vocabulary Semantic Segmentation

### v1

@@ Abstract
Open-vocabulary semantic segmentation aims to segment and recognize semantically meaningful regions based on text-based descriptions during inference. A typical solution to address this task is to leverage powerful vision-language models (VLMs), such as CLIP, to bridge the gap between open- and close-vocabulary recognition. As VLMs are usually pretrained with low-resolution images (e.g. $224\times 224$), most previous methods operate only on downscaled images. We question this design as low resolution features often fail to preserve fine details. Although employing additional image backbones for high-resolution inputs can mitigate this issue, it may also introduce significant computation overhead. Therefore, we propose MROVSeg, a multi-resolution training framework for open-vocabulary semantic segmentation with a single pretrained CLIP backbone, that uses sliding windows to slice the high-resolution input into uniform patches, each matching the input size of the well-trained image encoder. Its key components include a Multi-Res Adapter, which restores the spatial geometry and grasps local-global correspondences across patches by learnable convolutional and scale attention layers. To achieve accurate segmentation, we introduce Multi-grained Masked Attention scheme to aggregate multi-grained semantics by performing cross-attention between object queries and multi-resolution CLIP features within the region of interests. Through comprehensive experiments, we demonstrate the superiority of MROVSeg on well-established open-vocabulary semantic segmentation benchmarks, particularly for high-resolution inputs, establishing new standards for open-vocabulary semantic segmentation.
@@ 1 Introduction
(a) Previous open-vocabulary semantic segmentation.
(b) MROVSeg open-vocabulary semantic segmentation.
@@
Semantic segmentation is a fundamental computer vision problem which seek to group image pixels into semantically meaningful segments and recognize their categories. Previous research in semantic segmentation [25, 6, 9] mainly focus on segmenting objects from a pre-defined semantic categories set based on the dataset. These methods typically rely on human-annotated masks and class labels, which is costly and unscalable in terms to the number of categories. To this end, the task of open-vocabulary semantic segmentation has been introduced [39, 16, 21], which aims to segment semantic pixels belonging to arbitrary classes beyond pre-defined categories and dataset.
@@
In recent years, large-scale vision-language pretrained models (VLMs), such as CLIP [28] and ALIGN [19], have demonstrated remarkable generalization capabilities for recognizing open-vocabulary categories. This motivate the research community to investigate the potential of VLMs in open-vocabulary semantic segmentation. To address the discrepancy between the per-pixel semantic requirements and the image-level labels provided by VLMs, initial studies [11, 44, 45] modified CLIP model by removing its final pooling layer to obtain dense category embeddings for per-pixel classification. However, these approaches typically necessitate fine-tuning VLMs on a base segmentation dataset with limited images and categories, which is demonstrated [44] to impair the transferability of VLM features, leading to unsatisfactory zero-shot performance on downstream tasks.
@@
Recent approaches [35, 14, 36, 30, 24] reformulate open-vocabulary semantic segmentation as a region-level recognition problem. These methods typically adopt two branch meta architecture: one branch extract image feature and generate mask proposals, and the other branch classifies the predicted proposals with pretrained VLM. Although these methods are promising, we note their following limitations. First, due to pretrained VLMs exhibit inferior size adaptability, most of open-vocabulary semantic segmentation methods (e.g. [14, 36, 22, 35, 34, 8, 30]) so far need to downsample images to fit the pretrained resolution (e.g. $224\times 224$) of VLM to perform region-level recognition (as in Fig. 1 (a)). However, low-resolution input usually lacks segmentation details. Although naïvely applying sliding window inference [11, 35] could partly compensate for the details, the spatial structure across windows is corrupted and the local-global modeling is also absent.
@@
In light of the limitations and challenges faced by previous methods, we propose MROVSeg, a VLM-based Multi-Resolution training framework for Open-Vocabulary Semantic Segmentation. As illustrated in Fig.  1 (b), first, MROVSeg uses downsampled low-resolution images as VLM input to extract global low-resolution feature. Second, MROVSeg split the high-resolution images into slices and feeds them to VLM to extract detailed high-resolution features. The key components of MROVSeg contain a Multi-Res Adapter, in which we employs depthwise convolution layers to restore the spatial geometry across slices. To effectively capture global long-range context, inspired by previous multi-scale training framework [5, 18, 37], we employ a image-dependent Scale-aware Attention [5] to dynamically adjust the fusion weight of high-resolution and low-resolution VLM features based on their relevance and trustworthiness. The resulting multi-resolution features are fused hierarchically then employed for precise mask proposals generation.
@@
To achieve accurate mask class recognition, we propose a Multi-grained Masked Attention mechanism. The core insight is to reuse the CLIP [CLS] token, and manipulate its attention map on multi-resolution features in CLIP attention layers with resolution-aware attention masks. We find this resolution-aware design can enforce the low- and high-resolution attention map focus on global contexts and spatial details respectively and thus effectively aggregate multi-grained semantics.
With extensive experiments on well-established open-vocabulary semantic segmentation benchmarks, such as Pascal Context-59, Pascal Context-459 [26], ADE-150, ADE-847 [43], Pascal VOC [15], we are delighted to report that our method surpasses current state-of-the-art methods by a significant margin on standard benchmarks (+2.4 mIoU%), demonstrating the advancements of MROVSeg in the domain of open-vocabulary semantic segmentation. Our contributions can be summarized as follows:
@@
  • We propose a novel end-to-end multi-resolution training framework to tackle the task of open-vocabulary semantic segmentation. It enables improved open-vocabulary segmentation by leveraging multi-resolution vision-language features.
@@
  • The efficacy of our method is confirmed by achieving state-of-the-art performance by evaluating our proposed approach on five well-established open-vocabulary semantic segmentation benchmarks.
@@ 2 Related work
Figure 2: The overall pipeline of MROVSeg. For an high-resolution input image, its downsampled image and are fed into CLIP visual encoder to extract multi-resolution CLIP features. The multi-resolution CLIP features are injected into Multi-Res Adapter to produce multi-scale features for mask decoder.
Pretrained Vision-language Models. Recently, large-scale pretrained, contrastive learning based methods [28, 19] demonstrate powerful open-vocabulary image classification capability by learning shared image-text feature representations. Pretrained CLIP [28] has been generalized to many downstream computer vision tasks such as object detection [42], image caption [17], image generation [27] and image segmentation [36, 2, 34, 38, 35]. Our method invokes natural language perceptual ability of pretrained VLMs, aiming to explore their application boundary in open vocabulary semantic segmentation tasks.
Semantic Segmentation. Classic closed-vocabulary semantic segmentation interprets category concepts in a predefined label sets and generally performs segmentation on a specific dataset. In the past decade, FCN [25] and its variants [40, 6, 7, 41, 33] have dominated the field for several years by using per-pixel classification framework. Recently, MaskFormer [9, 10] explore to reformulate the semantic segmentation as a mask classification problem. Our approach is the extension of the traditional semantic segmentation methods, enabling them to segment a much broader range of semantic categories and adapt to real-world scenarios.
Open Vocabulary Image Segmentation. Open vocabulary image segmentation aims at segmenting and recognizing semantic regions described by text. Pioneering works [2, 31] use learned language embedding to align the feature space of class texts and visual semantics. Recently, some works [35, 14, 22, 24] develop two-stage training approaches, i.e. first train a stand-alone model to generate class-agnostic masks, and then adapt CLIP models to classify the mask proposals. More recently, some end-to-end frameworks emerge in the community, which unify mask generation and region classification into the same model. SAN [36] propose a lightweight side adapter to generate mask proposals and attention biases that act on CLIP attention layers. ODISE [34] employs a stable diffusion UNet [29] to generate mask proposals and adopt frozen CLIP model to enhance the open-vocabulary recognition capability. With the extracted dense pixel semantic embedding [44], CAT-Seg [11] proposes to finetune CLIP with cost aggregation. EBSeg [30] integrate powerful Segment Anything Model [20] into CLIP-based framework with image embedding balancing.
Discussion with Previous Methods. Our method MROVSeg is inspired by [36, 38], but has significant differences: (1) MROVSeg, which adapts multi-resolution CLIP feature to assist semantic segmentation by Multi-Res Adapter, is a multi-resolution training framework, and is different from all previous open-vocabulary semantic segmentation methods. (2) The introduced Muti-grained Masked Attention scheme explicitly enforces the mask class recognition to aggregate both local and global semantics, which takes advantage of the internal consistency between multi-resolution features.
@@ 3 Method
3.1 Problem Definition
In open vocabulary semantic segmentation task setting [35, 22], an image $I\in\mathbb{R}^{3\times H\times W}$ is operated by a segmentation model $\mathcal{P}_{\theta}$ with parameter $\theta$ to produce a set of masks associated with $K$ semantic categories:
@@ 3.1 Problem Definition
The segmentation model $\mathcal{P}_{\theta}$ is trained on a base segmentation dataset (e.g., COCO [3]) annotated with a fixed label set $\mathcal{Y}_{\mathrm{train}}=\{y_{i}\}_{i=1}^{K_{\mathrm{train}}}$ of $K_{\mathrm{train}}$ semantic categories. And during test, model $\mathcal{P}_{\theta}$ is expected to segment objects of category set $\mathcal{Y}_{\mathrm{test}}$, which generally contains novel categories ($\mathcal{Y}_{\mathrm{test}}\not\subset\mathcal{Y}_{\mathrm{train}}$).
3.2 Motivation
Accurate semantic segmentation needs high-resolution image inputs. Due to low-resolution images used in vision-language pretraining, previous open-vocabulary methods employ extra image backbone (such as SAM [30] and ResNet [24]) to provide segmentation detail. Although recent studies [32, 38] adapt convolution-based CLIP model for high-resolution training, but directly apply these methods to ViT-based CLIP model results in suboptimal performance [36] due to undesirable size adaptability of ViT. To this end, we propose MROVSeg, a ViT-based training framework to provide multi-resolution vision-language features for open-vocabulary semantic segmentation.
@@ 3.3 Framework Overview
The overview of our training framework MROVSeg, for open-vocabulary semantic segmentation is shown in Fig. 2. At a high-level, an image is downsampled and padded to a low-resolution (such as $224\times 224$) and processed by a pretrained CLIP ViT to extract global feature. To capture high-resolution local details, the high-resolution image is split into slices and input to the shared CLIP ViT encoder. These multi-resolution features are then concatenated with learnable queries and fed into a Multi-Res Adapter(Sec. 3.4), to produce the fused features, query features, and attention masks used for mask classification.
@@
3.4 Multi-Res Adapter
As depicted in Fig. 3, the Multi-Res Adapter adapts multi-resolution CLIP features for segmentation task. Denote the slice (high resolution) features of $l$-layer as $\{\mathbf{P}_{i}^{l}\}_{i=1}^{S}$, the global feature (low resolution) as $\bar{\mathbf{P}}_{i}^{l}$, where $\mathbf{P}_{i}\in\mathbb{R}^{L\times D}$, $S$ is the slice number, $L=H\times W$ is the token length, and $D$ is the channel number. Naïvely concatenating the high resolution slice features for the subsequent segmentation prediction is promising, but have two defects: 1) the spatial geometry, such as the positional information, is corrupted across slice features $\{\mathbf{P}_{i}^{l}\}_{i=1}^{S}$; 2) the long-range dependencies and global context is missing. Thus, we propose Multi-Res Adapter to effectively restore spatial geometry of slice features and capture long-range context from global feature. In Multi-Res Adapter, the 00-th slice features $\{\mathbf{P}_{i}^{0}\}_{i=1}^{S}$ are firstly concatenated with learnable queries $\mathbf{Q}\in\mathbb{R}^{N\times D}$ and input to the vanilla ViT blocks to build the query features for each objects. Then for target fusion layer $l$, the slice features $\{\mathbf{P}_{i}^{l}\}_{i=1}^{S}$ and global feature $\bar{\mathbf{P}}_{i}^{l}$ are fused through a Multi-Res Fusion (MRF) Module and then injected into the ViT branch.
@@ 3.6 Mask Classification
Decoupled Attention Mask Decoding. CLIP models, like general Vision Transformers, perform open-vocabulary image-level recognition through [CLS] token. Our goal is to recognize the various predicted mask proposals without changing any parameter of CLIP by guiding the attention maps of [CLS] token on the region of interests. To achieve this goal, we follow MaskCLIP [14] and SAN [36] to reuse the original CLIP [CLS] token in the intermediate layer. To sufficiently aggregate multi-grained semantics from CLIP, we first duplicate the [CLS] token to query number $N$ and create learnable positional embedding for them, dubbed as the $\mathrm{\mathbf{X}_{\texttt{prop}}}\in\mathbb{R}^{N\times D}$. We aim to enforce $\mathrm{\mathbf{X}_{\texttt{prop}}}$ to extract the global and local semantics from low- and high-resolution CLIP feature respectively. Thus, for visual feature $\mathbf{H}$, we first extract global contexts with max pooling $\bar{\mathbf{H}}=\texttt{MaxPool}(\mathbf{H})$ and train MLPs project them to attention space
@@
Multi-grained Masked Attention. As shown in Fig.4, we perform cross attention to update the $\mathrm{\mathbf{X}}_{\texttt{prop}}$ with multi-resolution CLIP features, with the predicted attention masks $\mathbf{M}_{\mathrm{local}}$ and $\mathbf{M}_{\mathrm{global}}$,
@@
Figure 4: Multi-grained Masked Attention.

### v2

# MROVSeg: Breaking the Resolution Curse of Vision-Language Models in Open-Vocabulary Image Segmentation

@@ Abstract
Pretrained vision-language models (VLMs), e.g. CLIP, are increasingly used to bridge the gap between open- and close-vocabulary recognition in open-vocabulary image segmentation. As VLMs are generally pretrained with low-resolution images (e.g. $224\times 224$), most previous methods operate only on downscaled images. We question this design as low resolution features often fail to preserve fine details. A typical solution is to employ additional image backbones for high-resolution inputs, but it also introduce significant computation overhead. Therefore, we propose MROVSeg, a multi-resolution training framework for open-vocabulary image segmentation with a single pretrained CLIP backbone, that uses sliding windows to slice the high-resolution input into uniform patches, each matching the input size of the well-trained image encoder. Its key components include a Multi-Res Adapter, which restores the spatial geometry and grasps local-global correspondences across patches by interacting with multi-resolution features. To achieve accurate segmentation, we introduce Multi-grained Masked Attention scheme to aggregate multi-grained semantics from multi-resolution CLIP features to object queries. Through comprehensive experiments, we demonstrate the superiority of MROVSeg on well-established open-vocabulary image segmentation benchmarks, establishing new standards for open-vocabulary image segmentation.
@@ 1 Introduction
(a) Previous open-vocabulary image segmentation.
(b) MROVSeg open-vocabulary image segmentation.
@@
Open-vocabulary image segmentation aims to segment semantic pixels belonging to arbitrary classes beyond pre-defined categories and dataset. In recent years, large-scale vision-language pretrained models (VLMs), such as CLIP [26] and ALIGN [14], have demonstrated remarkable generalization capabilities for recognizing open-vocabulary categories. This motivates the research community to investigate the potential of VLMs in open-vocabulary image segmentation. To address the discrepancy between the per-pixel semantic requirements and the image-level labels provided by VLMs, initial studies [6, 42, 43] modified CLIP model by removing its final pooling layer to obtain dense category embeddings for per-pixel classification. However, these approaches typically necessitate fine-tuning VLMs on a base segmentation dataset with limited images and categories, which is demonstrated [42] to impair the transferability of VLM features, leading to unsatisfactory zero-shot performance on downstream tasks.
@@
Recent approaches [35, 10, 36, 28, 22] reformulate open-vocabulary image segmentation as a region-level recognition problem. These methods typically adopt two branch meta architecture (As in Fig. 1(a)): one branch extract image feature and generate mask proposals, and the other branch classifies the predicted proposals with pretrained VLM. Although these methods are promising, we note their following limitation. Due to pretrained VLMs exhibit inferior size adaptability, most of open-vocabulary image segmentation methods (e.g. [10, 36, 19, 35, 34, 5, 28, 22]) so far need to downsample images to fit the pretrained resolution (e.g. $224\times 224$) of VLM to perform region-level recognition (as in Fig. 1(a)). However, low-resolution input usually lacks segmentation details. Although naïvely applying sliding window inference [6, 35] could partly compensate for the details, the spatial structure across windows is corrupted and the local-global modeling is also absent.
@@
In light of the limitations and challenges faced by previous methods, we propose MROVSeg, a VLM-based Multi-Resolution training framework for Open-Vocabulary Image Segmentation. As illustrated in  Fig. 1(b), first, MROVSeg uses downsampled low-resolution images as VLM input to extract global low-resolution features. Second, MROVSeg split the high-resolution images into slices and feeds them to VLM to extract detailed high-resolution features. The key components of MROVSeg contain a Multi-Res Adapter, in which we employs depthwise convolution layers to restore the spatial geometry across slices. To effectively capture global long-range context, inspired by previous multi-scale training framework [4, 13, 38], we employ a image-dependent Scale-aware Attention [4] to dynamically adjust the trustworthiness of high-resolution and low-resolution VLM features based on their relevance. The resulting multi-resolution features are fused hierarchically then employed for precise mask proposals generation.
@@
To achieve accurate mask class recognition, we propose a Multi-grained Masked Attention mechanism. The core hypothesis is that multi-resolution CLIP features of the save image input hold semantic consistency. Based on this, we reuse the CLIP [CLS] token, and manipulate its attention map on multi-resolution features in CLIP attention layers with resolution-aware attention masks. We find this resolution-aware design can enforce the low- and high-resolution attention map focus on global contexts and spatial details respectively and thus effectively aggregate multi-grained semantics. With extensive experiments on well-established open-vocabulary semantic segmentation and panoptic segmentation benchmarks, we are delighted to report that our method achieves new state-of-the-art performance, demonstrating the advancements of MROVSeg in the domain of open-vocabulary image segmentation. Our contributions can be summarized as follows:
@@
  • We propose a novel end-to-end multi-resolution training framework to tackle the task of open-vocabulary image segmentation. It enables improved open-vocabulary segmentation by leveraging multi-resolution vision-language features.
@@
  • The efficacy of our method is confirmed by achieving state-of-the-art performance by evaluating our proposed approach on well-established open-vocabulary segmentation benchmarks.
@@ 2 Related work
Figure 2: The overall pipeline of MROVSeg. For an high-resolution input image, its downsampled image and are fed into CLIP visual encoder to extract multi-resolution CLIP features. The Multi-Res Adapter adapts these features for mask decoder and attention mask decoder. The generated attention masks are employed to aggregate semantics from the multi-resolution CLIP features.
Pretrained Vision-language Models Recently, large-scale pretrained, contrastive learning based methods [26, 14] demonstrate powerful open-vocabulary image classification capability by learning shared image-text feature representations. Pretrained CLIP [26] has been generalized to many downstream computer vision tasks such as object detection [40], image caption [12], image generation [24] and image segmentation [36, 1, 34, 39, 35]. Our method invokes natural language perceptual ability of pretrained VLMs, aiming to explore their application boundary in open vocabulary semantic segmentation tasks.
Multi-Resolution Training As the quadratic computational overhead along with the number of tokens increases, recent multimodal large language models [21, 18, 37] (MLLMs) employ sliding window technique to divide high-resolution images into patches, thereby achieving competitive performance while maintaining computational efficiency. Unlike prevalent MLLMs, MROVSeg adaptively restores the spatial geometry of multi-resolution features across patches, and effectively extract global contexts that benefit segmentation tasks.
Open Vocabulary Image Segmentation Pioneering works [1, 32] use learned language embedding to align the feature space of class texts and visual semantics. Recently, some works [35, 10, 19, 22] develop two-stage training approaches. More recently, end-to-end [36] frameworks emerge in the community, which unify mask generation and region classification into the same model. SAN [36] propose a lightweight side adapter to effectively adapt CLIP features. ODISE [34] employs a stable diffusion UNet [27] to generate mask proposals. With the extracted dense pixel semantic embedding [42], CAT-Seg [6] proposes to finetune CLIP with cost aggregation. EBSeg [28] integrate Segment Anything Model [16] into CLIP-based framework with image embedding balancing.
Discussion with Previous Methods Our method MROVSeg is inspired by [36, 39], but has significant differences: (1) Distinct with previous open-vocabulary segmentation methods, MROVSeg adapts multi-resolution CLIP features to open-vocabulary segmentation. (2) The introduced Muti-grained Masked Attention scheme explicitly enforces the mask class recognition to aggregate both local and global semantics, which takes advantage of the internal consistency between multi-resolution features.
@@ 3 Method
In open vocabulary image segmentation task setting [35, 19], an image $I\in\mathbb{R}^{3\times H\times W}$ is operated by a segmentation model $\mathcal{P}_{\theta}$ with parameter $\theta$ to produce a set of masks associated with $K$ categories:
@@
The segmentation model $\mathcal{P}_{\theta}$ is trained on a base segmentation dataset (e.g., COCO [2]) annotated with a fixed label set $\mathcal{Y}_{\mathrm{train}}=\{y_{i}\}_{i=1}^{K_{\mathrm{train}}}$ of $K_{\mathrm{train}}$ categories. And during test, model $\mathcal{P}_{\theta}$ is expected to segment objects of category set $\mathcal{Y}_{\mathrm{test}}$, which generally contains novel categories ($\mathcal{Y}_{\mathrm{test}}\not\subset\mathcal{Y}_{\mathrm{train}}$).
Accurate segmentation needs high-resolution image inputs. Due to low-resolution images used in vision-language pretraining, previous open-vocabulary methods employ extra image backbone (such as SAM [28] and ResNet [22]) to provide segmentation detail. Although recent studies [33, 39] adapt convolution-based CLIP model for high-resolution training, but directly apply these methods to ViT-based CLIP model results in suboptimal performance [36] due to undesirable size adaptability of ViT. To this end, we propose MROVSeg, a ViT-based training framework to provide multi-resolution vision-language features for open-vocabulary image segmentation.
@@ 3.1 Framework Overview
The overview of our training framework MROVSeg, for open-vocabulary image segmentation is shown in Fig. 2. At a high-level, an image is downsampled and padded to a low-resolution (such as $224\times 224$) and processed by a pretrained CLIP ViT to extract global feature. To capture high-resolution local details, the high-resolution image is split into slices and input to the shared CLIP ViT encoder. These multi-resolution features are then concatenated with learnable queries and fed into a Multi-Res Adapter(Sec. 3.2), to produce the fused features, query features, and attention masks used for mask predction( Sec. 3.3) mask classification( Sec. 3.4).
3.2 Multi-Res Adapter
As depicted in Fig. 3, denote the slice (high resolution) features of $l$-layer as $\{\mathbf{P}_{i}^{l}\}_{i=1}^{S}$, the global feature (low resolution) as $\bar{\mathbf{P}}_{i}^{l}$, where $\mathbf{P}_{i}\in\mathbb{R}^{L\times D}$, $S$ is the slice number, $L=H\times W$ is the token length, and $D$ is the channel number. Naïvely concatenating the high resolution slice features for the subsequent segmentation prediction is promising, but have two defects: 1) the spatial geometry, such as the positional information, is corrupted across slice features $\{\mathbf{P}_{i}^{l}\}_{i=1}^{S}$; 2) the long-range dependencies and global context is missing. Thus, we propose Multi-Res Adapter to effectively restore spatial geometry of slice features and capture long-range context from global feature. In Multi-Res Adapter, the 00-th slice features $\{\mathbf{P}_{i}^{0}\}_{i=1}^{S}$ are firstly concatenated with learnable queries $\mathbf{Q}\in\mathbb{R}^{N\times D}$ and input to the vanilla ViT blocks to build the query features for each objects. Then for target fusion layer $l$, the slice features $\{\mathbf{P}_{i}^{l}\}_{i=1}^{S}$ and global feature $\bar{\mathbf{P}}_{i}^{l}$ are fused through a Multi-Res Fusion (MRF) Module and then injected into the ViT branch.
@@ 3.4 Mask Classification
Figure 4: Effect of decoupled attention decoding for multi-grained semantics. With single attention mask decoding, the spatial cues are overwhelmed by background noise (b). Our decoupled attention mask decoding effectively splits the global and local semantics, producing relatively clean global (c) and local (d) attention masks.
Recent ViT-based open-vocabulary segmentation methods [10, 36, 28] perform mask class recognition by predicting attention masks to guide the attention maps of original CLIP [CLS] token on the region of interests in the intermediate layers. We observe that the predicted attention masks (shown in Fig. 4(b)) in these methods tend to be overwhelmed by heavy background noises which spatially related to high norm tokens (shown in Fig. 4(a)), leading to the unsatisfactory classification performance. While prior ViT pretraining technique [8] indicates these high norm tokens contain rich global contexts, recent advances in training-free open-vocabulary segmentation [31, 29, 17] reveal that high norm tokens can easily disturb spatial correlation in CLIP features. Inspired by these inspection, we hypothesize that CLIP features hold internal consistency among multi-resolution inputs, and propose to simultaneously aggregate semantics from multi-resolution CLIP features by predicting decoupled attention masks.
Decoupled Attention Mask Decoding To sufficiently aggregate multi-grained semantics from CLIP, we first duplicate the [CLS] token to query number $N$ and create learnable positional embedding for them, dubbed as the $\mathrm{\mathbf{X}_{\texttt{prop}}}\in\mathbb{R}^{N\times D}$. We aim to enforce $\mathrm{\mathbf{X}_{\texttt{prop}}}$ to extract the global image-wise and local object-specific semantics from low- and high-resolution CLIP feature respectively. Thus, for visual feature $\mathbf{H}$, we first extract global contexts with max pooling $\bar{\mathbf{H}}=\texttt{MaxPool}(\mathbf{H})$ and train MLPs project them to attention space
@@
Multi-grained Masked Attention As shown in Fig.5, we perform cross attention to update the $\mathrm{\mathbf{X}}_{\texttt{prop}}$ with multi-resolution CLIP features, with the predicted attention masks $\mathbf{M}_{\mathrm{local}}$ and $\mathbf{M}_{\mathrm{global}}$,
@@
Figure 5: Multi-grained Masked Attention. Object [CLS] tokens $\mathrm{\mathbf{X}}_{\texttt{prop}}$ perform cross attention with high- and low-resolution CLIP features $\mathbf{X}_{\texttt{LR}}$ and $\mathbf{X}_{\texttt{HR}}$ with decoupled attention masks.
Image-conditioned Text Feature Recent studies [15, 6] reveal that CLIP text encoder struggles to generate discriminative text embeddings for similar categories. Thus, we follow MAFT-Plus [15] to condition the text embeddings with learnable cross attention layers between text embeddings and regional pooling visual features: $\mathbf{X}_{\texttt{text}}=\texttt{CrossAttnLayer}(\mathbf{X}_{\texttt{text}}^{ori},\texttt{MaxPool}(\mathbf{H}))$, where $\mathbf{H}$ is depicted in Sec. 3.2. $\mathbf{X}_{\texttt{text}}^{ori}$ is the original text embedding extract by CLIP text encoder.